Consider the following scenario to answer the following questions: On a particular Saturday, Mark Zuckerberg and Bill Gates can either plant trees or spread mulch in their gardens. Their maximum output per day is listed in the following table, along with spaces where you can calculate the opportunity cost.

Trees Planted Opportunity Cost of 1 Tree Amount of Mulch Spread (in cubic yards) Opportunity Cost of Spreading 1 Cubic
Yard of Mulch
Zuckerberg 20

30

Gates 15

30



At what terms of trade (relative price ratio) could they specialize and trade with one another so that both have more trees planted and mulch spread than they could accomplish on their own?

9 trees planted per 12 cubic yards of mulch spread

7 trees planted per 12 cubic yards of mulch spread

12 trees planted per 12 cubic yards of mulch spread

5 trees planted per 12 cubic yards of mulch spread

10 trees planted per 12 cubic yards of mulch spread
Answer #1

Option (2).

(a) For Zuckerberg,

Opportunity cost of 1 tree = 30/20 = 1.5 Mulch spread

Opportunity cost of 1 Mulch spread = 20/30 = 0.67 tree

(b) For Gates,

Opportunity cost of 1 tree = 30/15 = 2 Mulch spread

Opportunity cost of 1 Mulch spread = 15/30 = 0.5 tree

(c) Trade is beneficial when Terms of Trade (TOT) lies between both opportunity costs, so

0.5 tree < TOT of 1 Mulch spread < 0.67 Mulch spread.

(d) Among given options, only option (2) satisfies this condition since Opportunity cost of 1 Mulch spread = 7/12 = 0.58 &

0.5 < 0.58 < 0.67